OpenStudy (kl0723):

how do I do this? (-) integral of sinx(2x+2), that's a negative sign before the integral...

OpenStudy (anonymous):

\[-\int\sin x~(2x+2)~dx~~?\]

OpenStudy (kl0723):

yes... lol forgot to add the dx

OpenStudy (anonymous):

Integrate by parts by setting \[\begin{matrix} u=2x+2&&&dv=\sin x~dx\\ du=2~dx&&&v=-\cos x \end{matrix}\] \[\begin{align*}-\int\sin x~(2x+2)~dx&=-\left(-\cos x~(2x+2)-\int(-2\cos x)~dx\right)\\ &=(2x+2)\cos x+2\int\cos x~dx\\ &=(2x+2)\cos x+2\sin x+C \end{align*}\]

OpenStudy (kl0723):

Thank you SithsAndGiggles, this is actually a longer problem but for some reason I was getting stuck here because I didn't think I would need to integrate by parts twice in the same problem :)

OpenStudy (anonymous):

You're welcome!